# SeBS-Flow: Benchmarking Serverless Cloud Function Workflows

### v2

@@ IV-C Benchmark Suite
We follow standard design practices to build a new benchmark suite: it should be relevant, extensible, easy to use, and reproducible  [35, 36, 37, 11]. Our suite is relevant as we include applications representing a variety of workloads in the industry and academia (Section V). The implementation is based on an abstract workflow definition and can be extended to new platforms by implementing a single interface that transcribes our model definition to the new platform. To fulfill the two remaining criteria, we build our implementation upon SeBS [11], an established benchmark suite for FaaS: Benchmarks must be easy to deploy and execute to ensure their self-validation [35]. Integration into a maintained and up-to-date platform helps integrating new developments of serverless platforms continuously and avoids pushing this task to the end user. SeBS-Flow is multi-platform, supports automatic deployment of functions to the cloud, and integrates with services like storage and cloud logging, allowing developers to focus on the actual implementation rather than specifics of cloud providers, which can be time-consuming [38, 39].
@@
Serverless functions need cloud-managed storage to access data and retain state across invocations. To that end, SeBS automatically manages object storage instances and provides functions with a multi-cloud API. To create realistic workflow representations of web applications, we need to support low-latency data stores other than object storage. We chose NoSQL key-value storage for this task and extended SeBS with a high-level interface for creating, modifying, retrieving, and deleting items. The interface supports a partition and an optional sorting key. Each benchmark function can use multiple tables managed by the benchmark suite. We map the tables to DynamoDB on AWS, CosmosDB on Microsoft Azure, and Firestore in Datastore mode on Google Cloud.
@@
We collect timestamps for start and end of each function, its requestID, and a containerID to detect container reuse by using the temporary filesystem and global variables. The runtime of a phase is defined by the start of its earliest function and the end of the latest one. All collected values are sent to a Redis [40] instance deployed in the same cloud region. We chose an in-memory cache as it provides sub-millisecond latencies, reducing the risk of distorting the performance measurements.
V Benchmark Applications
@@
In SeBS-Flow, we implement six benchmarks covering real-life workloads. Also, we implement four microbenchmarks used in the evaluation: function chain, object storage performance, parallel invocations (Section VII-C1), and selfish detour (Section VII-C2). The selected benchmarks cover various domains that use workflows (Table IV), and correspond to previous findings on the characterization of workflow use cases [41, 10] regarding control-flow, number of functions, parallel invocations of the same functions, and longer runtimes: While 33% of workflows include complex control-flow, 50% are sequential, which we cover with Trip Booking and function chain microbenchmark. 72% of workflows use less than ten different functions, 52% involve parallel invocations of the same function, and 25% contain functions with a runtime of over a minute, which is also included in our benchmark suite. We visualize only one of the benchmarks here, but provide figures for the other benchmarks in the supplementary material.
@@
TABLE IV: Key features of different benchmarks.
@@
The benchmark detects objects in a video, and parallelizes the sequential benchmark in vSwarm [42] (Figure 6). Functions decode video frames and apply the Faster R-CNN model [43]. The decode function first downloads the video, decodes $F$ frames, and then uploads $N=\lceil\frac{F}{B}\rceil$ batches of size $B$. $N$ parallel detect functions compute $Y_{i}$, all detections with confidence $p>0.5$. Finally, detections are accumulated in acc, returning the final payload $Y$. We used $F=10$ frames and batch size $B=5$, yielding two parallel functions in the map phase.
Figure 6: The Video Analysis benchmark.
@@
The benchmark represents web applications, and it mocks a common example of reserving a hotel, car rental, and flight [44, 45]. The workflow is a pipeline of functions mocking the reservation system by storing trip data in a shared NoSQL database. It implements the SAGA pattern of long-running transactions [46] where a failure triggers the reversal of prior changes. For testing, we simulate failure in the last confirm function, which is followed by three consecutive functions to reverse the booking.
@@
We base our example on prior implementations [42, 47] and perform the standard problem of word counting. First, the split function partitions the input text into $N$ batches. $N$ parallel map functions count how often each word occurs in their text chunk next. Next, shuffle flattens the resulting array $Y_{i}|i<M$. Finally, $M$ reducers count the total occurrences of their respective word in parallel, yielding $Z_{i}$. The benchmark has two parameters: the number of mapping functions $N$, and the total number of words $W$. We set $N=3$ and $W=5000$, containing $M=5$ different words. MapReduce frameworks typically execute fully in parallel. However, the available workflow primitives necessitate the shuffle function, not relying on the array $Y_{i}$ itself but flattening it to enable the desired level of parallelism in reduce.
@@
ExCamera [48] uses interdependent video-processing tasks to encode videos in parallel. A video with $M$ total frames is processed in chunks of $N$ frames by $\frac{M}{N}=T$ parallel functions. First, each frame is encoded, yielding one key frame and $N-1$ interframes. Decode decodes all $N$ frames again, calculating the final state. The final state from the first frame of the chunk is used for reencoding the other frames, resulting in one final state and $N-2$ interframes. We derive our implementation from the original description of ExCamera [48] and the available implementation [49]. We use $M=30$ total frames and a chunk size of $N=6$, resulting in five parallel functions.
@@
This workload represents a typical training pipeline: It starts with gen generating a dataset, with the number of samples $N$ and the number of features $M$ as input. Then, we train $K$ different classifiers $C_{i}$ in parallel. We generate $N=500$ samples and $M=1024$ features, and train $K=2$ classifiers: a Support Vector Machine [50], and a Random Forest [51], creating two concurrent functions.
@@
This is a scientific workflow that identifies mutational overlaps using data from the 1000 Genomes project. It consists of five tasks and three phases: First, $N$ individuals functions parse the data for their chunk of the input file of size $M$ and then upload their results to the cloud storage. While individuals_merge merges the results to one, sifting computes the SIFT scores. In the last phase, mutation_overlap measures the overlap in SNP variants and frequency measures the frequency of mutation overlapping, both by population $P$. The benchmark has the number of lines as input $M$, number of parallel individuals functions $N$, and number of populations $P$ as input variables. We use $M=1250$ lines, $N=5$ parallel individuals function, and $P=6$ populations.
VI Evaluation of Workflow Model
@@
By reviewing existing literature on serverless workflows, we evaluate whether our model is general enough to express applications of workflows and if our transcription to the platform-specific representations adds overhead compared to the native implementation. We do so by using the meta-search engine Google Scholar to find peer-reviewed publications containing the keywords cloud, orchestration, and serverless workflow or serverless DAG. We exclude papers that are not in English, do not use a workflow benchmark, or are published before 2017, the year of the first serverless workflows in the cloud. This results in 72 papers analyzed papers (cf. Table I, p. I for their categorization). We provide the complete list of papers and analysis results in the supplementary material.
VI-A Expressiveness of our Model
@@
We analyze the workflow benchmarks used in the literature and evaluate whether our model can represent the control flow within the workflows without adding unnecessary dependencies between their tasks. Out of the 72 papers, 14 did not provide sufficient detail on the workflows used and their dependencies to judge if we can express them. In two papers, benchmarks are not presentable by our model, as they introduce new programming models to support communication between functions and load-balanced orchestration. Benchmarks used in three more papers can be modeled but not transcribed to platform-specific representations (Section IV). For two of them, cloud platforms are the limitations, such as ending the workflow as a result of a switch state (not possible on AWS) and using multi-stage inputs, i.e., using the output of a previously executed function as input without passing it to the functions invoked in-between. While we do not currently support transcribing the switch state requiring two conditions to be true, it can be easily added to the implementation. We fully support modeling and transcribing the workflows described in 53 of the 58 analyzed papers. Therefore, we conclude that our model does not have general limitations and developers can use it to model and execute their workflows.
VI-B Overhead of our Model
@@
To check if our model and transcription create overhead compared to a native implementation, we evaluate available benchmark implementations used in the analyzed papers and compare them to our transcription of their workflows. Only 10 of the 72 papers include an artifact containing workflow implementations or show their implementation as part of the paper for any of the platforms we support. None of them uses Google Cloud Workflows. In total, we find eleven AWS Step Functions state machines. One of them uses the AND choice type we do not transcribe, and another one adds fail and success states before ending the workflow, which only introduces overhead as compared to just ending the workflow. The other nine state machines use the same states with the same parameters in the same order as the state machines we transcribe, except for the fact that they specify each parameter explicitly as part of the state machine while we wrap them within a single payload entry. Four of the papers provide implementations for a total of six workflows using Azure Durable Functions. While one paper only provides an implementation using entity functions, the other five workflow implementations use activities to orchestrate tasks similar to our transcription. Since we must parse the platform-independent representation within the orchestrator, we could introduce an overhead. However, the evaluation of the 1000Genome benchmark, the benchmark with the most functions, shows that the average duration of the orchestrator function is only 13.6 milliseconds. We conclude that SeBS-Flow does not introduce noteworthy overhead in the workflows compared to their native implementation, enabling developers to obtain realistic performance results for their workflows.
VI-C Threats to Validity
@@
We used only one query to find relevant works, bearing the risk of missing results. We mitigated this by evaluating different queries beforehand, evaluating the relevance of papers found and checking if relevant papers we knew were included. Regarding external validity, we found only a limited number of artifacts to evaluate the overhead, with none available that uses GC Workflows. While our transcription follows best practices and tutorials as provided by the cloud providers and matches the artifacts we found, usage in other projects could differ.
VII Evaluation of Cloud Services
@@
We use SeBS-Flow to evaluate three major cloud workflow services – AWS Step Functions, Google Cloud Workflows, and Azure –providing developers valuable insights regarding their suitability for different workloads. We investigate the following research questions:
@@ 2. RQ2
    2. RQ2.2
      What causes overheads between function invocations?
@@ 5. RQ5
VII-A Methodology
We deploy benchmarks on Azure to the europe-west region, on AWS to us-east-1, and on Google Cloud to us-east1. We use the lowest common memory configuration that successfully executes the workflow on AWS and Google Cloud, at least 256 MB for computational functions and 128 MB for simple web applications. We invoke the application benchmarks in burst mode, triggering 30 executions at once and accepting all successful workflow executions, as other work suggests that most serverless applications have potentially bursty workloads [10]. We check how often we should repeat experiments by computing non-parametric confidence intervals on the results for the MapReduce benchmark and aim at being in a 5% interval of the median with a 95% confidence interval. For the burst mode with 30 executions triggered at once, this results in 1, 1, and 6 repetitions on AWS, GCP, and Azure, respectively. We opt to execute all experiments 180 times. However, we could only obtain 30 executions of the 1000Genome benchmark on Azure due to frequent timeout issues. Benchmarks use the serverless object storage and NoSQL database on each platform.
VII-B RQ1: Runtime Differences among Platforms
We compare the runtime of each benchmark on selected platforms. We calculate the runtime by subtracting the first start timestamp from the last end timestamp. The results in Figure 7 do not yield a single fastest platform among all our benchmarks. AWS is the fastest platform for three out of six benchmarks while performing relatively well for the other three. While Google Cloud’s performance is comparable to AWS, it is 1.55-1.97x slower on three benchmarks. While Azure Durable functions perform very well, e.g., on MapReduce and Machine Learning, they are the slowest platform for Video Analysis, ExCamera, and the 1000Genome benchmark. For Trip Booking, Azure achieves the best median performance but suffers from large outliers. We investigate the potential causes of slowdown in the next section. All platforms demonstrate variable performance, with Azure showing the largest variance.

### v3

@@ 4.3. Benchmark Suite
We follow standard design practices to build a new benchmark suite: it should be relevant, extensible, easy to use, and reproducible  (v. Kistowski et al., 2015; Binnig et al., 2009; Hoefler and Belli, 2015; Copik et al., 2021). Our suite is relevant as we include applications representing a variety of workloads in the industry and academia (Section 5). The implementation is based on an abstract workflow definition and can be extended to new platforms by implementing a single interface that transcribes our model definition to the new platform. To fulfill the two remaining criteria, we build our implementation upon SeBS (Copik et al., 2021), an established benchmark suite for FaaS: Benchmarks must be easy to deploy and execute to ensure their self-validation (v. Kistowski et al., 2015). Integration into a maintained and up-to-date platform helps integrating new developments of serverless platforms continuously and avoids pushing this task to the end user. SeBS-Flow is multi-platform, supports automatic deployment of functions to the cloud, and integrates with services like storage and cloud logging, allowing developers to focus on the actual implementation rather than specifics of cloud providers, which can be time-consuming (Chatley and Allerton, 2020; Ristov et al., 2024).
@@
Serverless functions need cloud storage to access data and retain state across invocations. To that end, SeBS automatically manages object storage instances and provides functions with a multi-cloud API. To create realistic workflow representations of web applications, we need to support low-latency data stores other than object storage. We chose NoSQL key-value storage for this task and extended SeBS with a high-level interface for creating, modifying, retrieving, and deleting items. The interface supports a partition and an optional sorting key. Each benchmark function can use multiple tables managed by the benchmark suite. We map the tables to DynamoDB on AWS, CosmosDB on Microsoft Azure, and Firestore in Datastore mode on Google Cloud.
@@
We collect timestamps for start and end of each function, its requestID, and a containerID to detect container reuse by using the temporary filesystem and global variables. The runtime of a phase is defined by the start of its earliest function and the end of the latest one. All collected values are sent to a Redis (Redis, 2024) instance deployed in the same cloud region. We chose an in-memory cache as it provides sub-millisecond latencies, reducing the risk of distorting the performance measurements.
To implement a workflow using our model, a user has to provide the following (cf. Figure 6): First, the implementation of the workflow’s functions in a language of their choice. Our workflow model is independent of the actual benchmark implementation. We can work with any language supported in the cloud, with currently supported Python, Node.js, C++, and Java through SeBS. Second, any data used as input to the workflow. Third, the specification how the functions should be orchestrated using our platform-agnostic workflow definition in JSON (cf. subsection 4.1). SeBS-Flow takes both as input and deploys the workflow with the functions to the respective cloud the user chooses, transcribing the workflow to their platform-specific representations by traversing the JSON file. This is transparent to the user and fully automated.
Figure 6. Process of executing a workflow using SeBS-Flow.
5. Benchmark Applications
@@
In SeBS-Flow, we implement six benchmarks covering real-life workloads. Also, we implement four microbenchmarks used in the evaluation: function chain, object storage performance, parallel invocations (Section 7.3.1), and selfish detour (Section 7.3.2). The selected benchmarks cover various domains that use workflows (Table 4), and correspond to previous findings on the characterization of workflow use cases (Eismann et al., 2020, 2022) regarding control-flow, number of functions, parallel invocations of the same functions, longer runtimes, and workload sizes: We include the sequential TripBooking benchmark (50% of workflows), four benchmarks using less than ten different functions (72% of workflows), five benchmarks involving parallel invocations of the same function (52% of workflows), and the 1000Genome workflow contains functions with a runtime of over a minute (25% of workflows). Moreover, the analysis by Eismann et al. (2022) classifies 72% of serverless workflows as “small,” consisting of 2 to 10 functions, 23% as “medium” with 11 to 1000 functions, and only 4% as “large” with more than 1000 functions. With our workloads, four of six applications can be classified as “small” and two more as “medium.” Furthermore, the analysis of traces from production workloads on Azure Durable Functions (Mahgoub et al., 2022) also showed that 40% of workflows are sequential, most workflows do not use more than ten functions, the median number of different functions in a workflow is three, and the median workflow execution time is 5.6 seconds, confirming the representativeness of our applications. We visualize only one of the benchmarks here, but provide figures for the other benchmarks in the supplementary material.
@@
Table 4. Key features of different benchmarks.
@@
The benchmark detects objects in a video, and parallelizes the sequential benchmark in vSwarm (vsw, 2024) (Figure 7). Functions decode video frames and apply the Faster R-CNN model (Ren et al., 2015). The decode function first downloads the video, decodes $F$ frames, and then uploads $N=\lceil\frac{F}{B}\rceil$ batches of size $B$. $N$ parallel detect functions compute $Y_{i}$, all detections with confidence $p>0.5$. Finally, detections are accumulated in acc, returning the final payload $Y$. We used $F=10$ frames and batch size $B=5$, yielding two parallel functions in the map phase.
Figure 7. The Video Analysis benchmark.
@@
The benchmark represents web applications, and it mocks a common example of reserving a hotel, car rental, and flight (McCaffrey, 2015; sag, 2023). The workflow is a pipeline of functions mocking the reservation system by storing trip data in a shared NoSQL database. It implements the SAGA pattern of long-running transactions (Garcia-Molina and Salem, 1987) where a failure triggers the reversal of prior changes. For testing, we simulate failure in the last confirm function, which is followed by three consecutive functions to reverse the booking.
@@
We base our example on prior implementations (vsw, 2024; MapReduce, 2024) and perform the standard problem of word counting. First, the split function partitions the input text into $N$ batches. $N$ parallel map functions count how often each word occurs in their text chunk next. Next, shuffle flattens the resulting array $Y_{i}|i<M$. Finally, $M$ reducers count the total occurrences of their respective word in parallel, yielding $Z_{i}$. The benchmark has two parameters: the number of mapping functions $N$, and the total number of words $W$. We set $N=3$ and $W=5000$, containing $M=5$ different words. MapReduce frameworks typically execute fully in parallel. However, the available workflow primitives necessitate the shuffle function, not relying on the array $Y_{i}$ itself but flattening it to enable the desired level of parallelism in reduce.
@@
ExCamera (Fouladi et al., 2017) uses interdependent video-processing tasks to encode videos in parallel. A video with $M$ total frames is processed in chunks of $N$ frames by $\frac{M}{N}=T$ parallel functions. First, each frame is encoded, yielding one key frame and $N-1$ interframes. Decode decodes all $N$ frames again, calculating the final state. The final state from the first frame of the chunk is used for reencoding the other frames, resulting in one final state and $N-2$ interframes. We derive our implementation from the original description of ExCamera (Fouladi et al., 2017) and the available implementation (Fouladi et al., 2019). We use $M=30$ total frames and a chunk size of $N=6$, resulting in five parallel functions.
@@
This workload represents a typical training pipeline: It starts with gen generating a dataset, with the number of samples $N$ and the number of features $M$ as input. Then, we train $K$ different classifiers $C_{i}$ in parallel. We generate $N=500$ samples and $M=1024$ features, and train $K=2$ classifiers: a Support Vector Machine (Platt, 1999), and a Random Forest (Breiman, 2001), creating two concurrent functions.
1000Genomes
@@
This is a scientific workflow that identifies mutational overlaps using data from the 1000 Genomes project (100, 2025). It consists of five tasks and three phases: First, $N$ individuals functions parse the data for their chunk of the input file of size $M$ and then upload their results to the cloud storage. While individuals_merge merges the results to one, sifting computes the Sorting Intolerant from Tolerant (SIFT) scores. In the last phase, mutation_overlap measures the overlap in Single Nucleotide Polymorphisms (SNP) variants and frequency measures the frequency of mutation overlapping, both by population $P$. The benchmark has the number of lines as input $M$, number of parallel individuals functions $N$, and number of populations $P$ as input variables. We use $M=1250$ lines, $N=5$ parallel individuals function, and $P=6$ populations.
6. Evaluation of Workflow Model
@@
By reviewing existing literature on serverless workflows, we evaluate whether our model is general enough to express applications of workflows and if our transcription to the platform-specific representations adds overhead compared to the native implementation. We do so by using the meta-search engine Google Scholar to find peer-reviewed publications containing the keywords cloud, orchestration, and serverless workflow or serverless DAG. We exclude papers that are not in English, do not use a workflow benchmark, or are published before 2017, the year of the first serverless workflows in the cloud. This results in 72 papers analyzed papers (cf. Table 1, p. 1 for their categorization). We provide the complete list of papers and analysis results in the supplementary material.
6.1. Expressiveness of our Model
@@
We analyze the workflow benchmarks used in the literature and evaluate whether our model can represent the control flow within the workflows without adding unnecessary dependencies between their tasks. Out of the 72 papers, 14 did not provide sufficient detail on the workflows used and their dependencies to judge if we can express them. In two papers, benchmarks are not presentable by our model, as they introduce new programming models to support communication between functions and load-balanced orchestration. Benchmarks used in three more papers can be modeled but not transcribed to platform-specific representations (Section 4). For two of them, cloud platforms are the limitations, such as ending the workflow as a result of a switch state (not possible on AWS) and using multi-stage inputs, i.e., using the output of a previously executed function as input without passing it to the functions invoked in-between. The third one uses a switch state requiring two conditions to be true. While we do not support transcribing this currently, transcription the switch state requiring two conditions to be true, it can be easily added to the implementation. We fully support modeling and transcribing the workflows described in 53 of the 58 analyzed papers. Therefore, we conclude that our model does not have general limitations within the scope of programming models not allowing for communication between functions and using orchestration based on dynamic characteristics of the system, and developers can use it to model and execute their workflows.
6.2. Overhead of our Model
@@
To check if our model and transcription (cf. Sec. 4.2) create overhead compared to a native implementation, we evaluate available benchmark implementations used in the analyzed papers and compare them to our transcription of their workflows. Only 10 of the 72 papers include an artifact containing workflow implementations or show their implementation as part of the paper for any of the platforms we support. None of them uses Google Cloud Workflows. In total, we find eleven AWS Step Functions state machines. One of them uses the AND choice type. We currently do not transcribe this choice type and are therefore not able to generate the same state machine. However, if we would add the transcription, the resulting state machine would look similar. Another one adds fail and success states before ending the workflow, which only introduces overhead as compared to just ending the workflow. The other nine state machines use the same states with the same parameters in the same order as the state machines we transcribe, except for the fact that they specify each parameter explicitly as part of the state machine while we wrap them within a single payload entry, which does not affect the overhead. Four of the papers provide implementations for a total of six workflows using Azure Durable Functions. While one paper only provides an implementation using entity functions, the other five workflow implementations use activities to orchestrate tasks similar to our transcription. Since we must parse the platform-independent representation within the orchestrator, we could introduce an overhead. However, the evaluation of the 1000Genome benchmark, the benchmark with the most functions, shows that the average duration of the orchestrator function is only 13.6 milliseconds, with the workflow’s median runtime being 3757.55 seconds. We conclude that SeBS-Flow does not introduce noteworthy overhead in the workflows compared to their native implementation, enabling developers to obtain realistic performance results for their workflows.
6.3. Threats to Validity
@@
We used only one query to find relevant works, bearing the risk of missing results. We mitigated this by evaluating different queries beforehand, evaluating the relevance of papers found, and checking if the results included relevant papers we knew as a gold standard (Dieste et al., 2009). Regarding external validity, we found only a limited number of artifacts to evaluate the overhead, with none available that uses GC Workflows. While our transcription follows best practices and tutorials as provided by the cloud providers and matches the artifacts we found, usage in other projects could differ.
7. Evaluation of Cloud Services
@@
We use SeBS-Flow to evaluate three major cloud workflow services – AWS Step Functions, Google Cloud Workflows, and Azure Durable Functions – providing developers valuable insights regarding their suitability for different workloads. We investigate the following research questions:
@@ 5. RQ5
7.1. Methodology
We deploy benchmarks and resources used by them on Azure to the europe-west region, on AWS to us-east-1, and on Google Cloud to us-east1. We use the lowest common memory configuration that successfully executes the workflow on AWS and Google Cloud, at least 256 MB for computational functions and 128 MB for simple web applications. We invoke the application benchmarks in burst mode, triggering 30 executions at once and accepting all successful workflow executions, as other work suggests that most serverless applications have potentially bursty workloads (Eismann et al., 2022). We check how often we should repeat experiments by computing non-parametric confidence intervals on the measurements of the MapReduce benchmark and aim at being in a 5% interval of the median using a 95% confidence interval. For the burst mode with 30 executions triggered at once, this results in 1, 1, and 6 repetitions on AWS, GCP, and Azure, respectively. We opt to execute all experiments 180 times. However, we could only obtain 30 executions of the 1000Genome benchmark on Azure due to frequent timeout issues. Benchmarks use the serverless object storage and NoSQL database on each platform.
7.2. RQ1: Runtime Differences among Platforms
We compare the runtime of each benchmark on the selected platforms. We calculate the runtime by subtracting the first start timestamp from the last end timestamp. The results in Figure 8 do not yield a single fastest platform among all our benchmarks. AWS is the fastest platform for three out of six benchmarks while performing relatively well for the other three. While Google Cloud’s performance is comparable to AWS, it is 1.55-1.97x slower on three benchmarks. While Azure Durable functions perform very well, e.g., on MapReduce and Machine Learning, they are the slowest platform for Video Analysis, ExCamera, and the 1000Genome benchmark. For Trip Booking, Azure achieves the best median performance but suffers from large outliers. We investigate the potential causes of slowdown in the next section. All platforms demonstrate variable performance, with Azure showing the largest variance.